### v2

# Diff-Instruct*: Towards Human-Preferred One-step Text-to-image Generative Models

Abstract
In this paper, we introduce the Diff-Instruct* (DI*), an image data-free approach for building one-step text-to-image generative models that align with human preference while maintaining the ability to generate highly realistic images. We frame human preference alignment as online reinforcement learning using human feedback (RLHF), where the goal is to maximize the reward function while regularizing the generator distribution to remain close to a reference diffusion process. Unlike traditional RLHF approaches, which rely on the KL divergence for regularization, we introduce a novel score-based divergence regularization, which leads to significantly better performances. Although the direct calculation of this preference alignment objective remains intractable, we demonstrate that we can efficiently compute its gradient by deriving an equivalent yet tractable loss function. Remarkably, we used Diff-Instruct* to train a Stable Diffusion-XL-based 1-step model, the 2.6B DI*-SDXL-1step text-to-image model, which can generate images of a resolution of $1024\times 1024$ with only 1 generation step. DI*-SDXL-1step model uses only 1.88% inference time and 29.30% GPU memory cost to outperform 12B FLUX-dev-50step significantly in PickScore, ImageReward, and CLIPScore on Parti prompt benchmark and HPSv2.1 on Human Preference Score benchmark, establishing a new state-of-the-art benchmark of human-preferred 1-step text-to-image generative models. Besides the strong quantitative performances, extensive qualitative comparisons also confirm the advantages of DI* in terms of maintaining diversity, improving image layouts, and enhancing aesthetic colors. We have released our industry-ready model on the homepage: https://github.com/pkulwj1994/diff_instruct_star.
1 Introductions
Generative models have made substantial progress in recent years, largely improving the content creation and editing across various domains (Karras et al., 2020; Nichol & Dhariwal, 2021; Poole et al., 2022; Kim et al., 2022; Tashiro et al., 2021; Meng et al., 2021; Couairon et al., 2022; Ramesh et al., 2022; Esser et al., 2024; Oord et al., 2016; Ho et al., 2022; Brooks et al., 2024; Zhang et al., 2023a; Xue et al., 2023; Luo & Zhang, 2024; Luo et al., 2023b; Zhang et al., 2023b; Feng et al., 2023; Deng et al., 2024; Luo et al., 2024d; Geng et al., 2024; Wang et al., 2024; Pokle et al., 2022).
Particularly, two types of generative models have gained significant attention, diffusion models and one-step generators. Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020), or score-based generative models (Song et al., 2020), progressively corrupt data with diffusion processes and then train models to approximate the score functions of the noisy data distributions across varying noise levels. The learned score functions can generate high-quality samples by denoising the noisy samples through reversed stochastic differential equations. While DMs can produce good outputs, they often require a large number of model evaluations, which limits their efficiency in applications.
Different from diffusion models, one-step generators (Luo et al., 2024c; a; Zheng & Yang, 2024; Kang et al., 2023; Sauer et al., 2023a; Yin et al., 2024; Zhou et al., 2024a) have emerged as a highly efficient alternative to multi-step diffusion models. Unlike DMs, one-step generative models directly map latent noises as inputs to generate data samples with a single neural network forward pass. This 1-step generation mechanism makes them ideal for real-time applications such as text-to-image and text-to-video generations on edge devices. Many existing works have demonstrated the leading performances of one-step text-to-image generators (Luo et al., 2024c; Zhou et al., 2024a; Yin et al., 2024) by employing diffusion distillation (Luo, 2023) as well as GAN techniques(Goodfellow et al., 2014; Brock et al., 2018; Karras et al., 2020; Sauer et al., 2023b). However, these works only focus on matching the 1-step model’s distributions with pre-trained diffusion models or ground truth data, without considering the critical challenge of teaching one-step text-to-image models to satisfy human preferences, which is one of the most important needs in the age of human-centric AI.
Figure 1: Generated $1024\times 1024$ images from 1-step 2.6B DI*-SDXL model. The DI*-SDXL-1step model outperforms FLUX-dev-50step(flu, ) and SD3.5-large-28step(sd, 3) in human preference metrics on the Parti prompt benchmark and HPSv2.1 scores. Please zoom in to check the details.
To resolve this issue, we introduce Diff-Instruct* (DI*), a novel approach that can train one-step text-to-image generators to have strong human-preference alignment while still maintaining the ability to generate diverse and realistic high-resolution images. We frame the human-preference alignment problem as maximizing the rewards with a score-based divergence constraint. This yields generated samples that not only have improved human reward but also adhere to user prompts. Different from traditional RLHF methods (Christiano et al., 2017; Ouyang et al., 2022; Luo, 2024), which rely on Kullback-Leibler (KL) divergence for distribution regularization, we find score-based regularization maintains better diversities, preventing an aligned 1-step model from stupidly generating painting-like images that have high human reward scores but poor diversity.
In our experiments, we use Diff-Instruct* to train 1-step text-to-image models based on SDXL(Podell et al., 2023) reference diffusion. We introduced the DI*-SDXL-1step model, an industry-ready model with 2.6B parameters and a UNet architecture. We evaluate our DI*-SDXL-1step model with other leading models extensively on both Parti prompts (Yu et al., 2022), COCO prompts, and Human Preference Score (HPS) v2.1 benchmarks. On Parti-prompt benchmark, the DI*-SDXL-1step model shows a very strong performance, outperforming 12B FLUX-dev-50step on PickScore, ImageReward, and CLIPScore, with only 1.88% inference time and 29.30% GPU memory cost. On COCO prompt datasets, the DI*-SDXL-1step goes on par with FLUX-dev and SD3.5-large diffusion models. On the Human Preference Score benchmarks, the 2.6B DI*-SDXL-1step model achieves a SoTA overall HPSv2.1 score of 31.19, demonstrating very strong human preference performances among all open-sourced models including Diffusio-DPO, DMD2, and FLUX-dev models. These solid results establish the Diff-Instruct* as a simple yet powerful approach for preference alignment of text-to-image models with SoTA open-sourced 1-step models.
We summarize our contribution in the following points:
  • We introduced Diff-Instruct* a novel approach for preference alignment of 1-step generative models with theoretical guarantees based on a novel score-based PPO paradigm;
  • We introduce two novel approaches to incorporate traditional classifier-free guidance into human preference alignment through explicit-implicit reward decoupling.
  • With extensive evaluations on three standard benchmarks across different alignment configurations, our best open-sourced 2.6B DI*-SDXL-1step text-to-image model, a leading text-to-image model with a resolution of $1024\times 1024$ that significantly outperforms previous leading FLUX-dev-50step model with only 1.88% inference time.
2 Preliminary
Diffusion Models.
Here we introduce preliminary knowledge and notations about diffusion models. Let $q_{0}(\bm{x})=q_{d}(\bm{x})$ be the data distribution. The goal of generative modeling is to train models to generate new samples $\bm{x}\sim q_{0}(\bm{x})$. Under mild conditions, the forward diffusion process of a diffusion model can transform initial distribution $q_{0}$ towards some simple noise distribution,
$\mathrm{d}\bm{x}_{t}=\bm{F}(\bm{x}_{t},t)\mathrm{d}t+G(t)\mathrm{d}\bm{w}_{t},$ (2.1)
where $\bm{F}$ is a pre-defined vector-valued drift function, $G(t)$ is a pre-defined scalar-value diffusion coefficient, and $\bm{w}_{t}$ denotes an independent Wiener process. A continuous-indexed score network $\bm{s}_{\varphi}(\bm{x},t)$ is employed to approximate marginal score functions of the forward diffusion process (2.1). The learning of score networks is achieved by minimizing a weighted denoising score matching objective (Vincent, 2011; Song et al., 2020),
$\mathcal{L}_{DSM}(\varphi)=\int_{t=0}^{T}\lambda(t)\mathbb{E}_{\bm{x}_{0}\sim q_{0},\atop\bm{x}_{t}|\bm{x}_{0}\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})}\|\bm{s}_{\varphi}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\|_{2}^{2}\mathrm{d}t.$ (2.2)
The weighting function $\lambda(t)$ controls the importance of the learning at different time levels and $p_{t}(\bm{x}_{t}|\bm{x}_{0})$ denotes the conditional transition of the forward diffusion (2.1). After training, $\bm{s}_{\varphi}(\bm{x}_{t},t)\approx\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})$ is a good approximation of marginal scores of diffused data distribution.
3 Preference Alignment using Diff-Instruct*
Problem Setup.
Our basic setting is that we have a known human reward function $r(\bm{x}_{0},\bm{c})$, which encodes the human preference for an image $\bm{x}_{0}$ and corresponding text description $\bm{c}$. Besides, we also have a pre-trained diffusion model which will later act as a reference distribution $p_{ref}(\bm{x}_{0})=q_{0}(\bm{x}_{0})$. The reference diffusion model is specified by the score function $\bm{s}_{q_{t}}(\bm{x}_{t})\coloneqq\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})$, where $q_{t}(\bm{x}_{t})$’s is the underlying distribution diffused at time $t$ according to (2.1). The pre-trained diffusion model represents the ground-truth data distribution that is used to prevent the aligned 1-step model from seeking high human reward but losing the ability to generate realistic images.
Our goal is to train a human-preferred one-step text-to-image model $g_{\theta}$ that generates images by directly mapping a random noise $\bm{z}\sim p_{z}$ to obtain $\bm{x}_{0}=g_{\theta}(\bm{z},\bm{c})$, conditioned on the input text $\bm{c}\sim\mathcal{C}$. We want the 1-step model’s output distribution, $p_{\theta}(\bm{x}_{0}|\bm{c})$, to maximize the expected human rewards while maintaining the ability to generate realistic images. We force $p_{\theta}(\bm{x}_{0}|\bm{c})$ to not diverge from $p_{ref}(\cdot)$ by using a divergence regularization term. Let $\bm{D}(\cdot,\cdot)$ be a distribution divergence. For any fixed prompt $\bm{c}$, the training objective is defined as:
$\theta^{*}=\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{\bm{x}_{0}\sim p_{\theta}(\bm{x}_{0}|\bm{c})}\big{[}-\alpha r(\bm{x}_{0},\bm{c})\big{]}+\bm{D}(p_{\theta},p_{ref})$ (3.1)
Here $\alpha$ is a coefficient that balances reward influences and $\bm{D}(\cdot)$ acts as a regularization term. When using Kullback-Leibler divergence as the regularization term, the objective (3.1) turns to the online Proximal Policy Gradient (PPO) algorithm(Schulman et al., 2017), which has great success in reinforcement learning from human feedback (RLHF) for large language models (Ouyang et al., 2022) and one-step text-to-image models alignment (Luo, 2024). Particularly, recently (Luo, 2024) has shown that traditional KL-PPO can fairly train human-preferred one-step text-to-image models. However, KL divergence is notorious for mode-seeking behavior, which makes Luo (2024) very easy to collapse to some distribution modes, such as painting-like images, that have very high human reward but lack diversity. To address this issue, we find that general score-based divergences have an appealing diversity-preserving property that shows great successes in 1-step diffusion distillation (Luo et al., 2024c; Zhou et al., 2024b). Based on such observations, we propose Diff-Instruct*, a novel practical algorithm based on an innovative Score-based online PPO paradigm that uses a general score-based divergence as regularization for reinforcement learning using human feedback.
Figure 2: Overall workflow of developing 1-step text-to-image generative models using Diff-Instruct*. The workflow includes stages of pre-training, reward modeling, and preference alignment.
3.1 General Score-based Divergences
Inspired by score-implicit matching (Luo et al., 2024c) that revealed the advantages of score-based divergences over KL divergences, we define the regularization term $\bm{D}(p_{\theta},p_{ref})$ via the following general score-based divergence. Assume $\mathbf{d}:\mathbb{R}^{d}\to\mathbb{R}$ is a scalar-valued proper distance function (i.e., a non-negative function that satisfies $\forall\bm{x},\mathbf{d}(\bm{x})\geq 0$ and $\mathbf{d}(\bm{x})=0$ if and only if $\bm{x}=\bm{0}$). Given a parameter-independent sampling distribution $\pi_{t}$ that has large distribution support, we can formally define a time-integral score divergence as
$\mathbf{D}^{[0,T]}(p_{\theta},p_{ref})\coloneqq\int_{t=0}^{T}w(t)\mathbb{E}_{\bm{x}_{t}\sim\pi_{t}}\bigg{\{}\mathbf{d}(\bm{s}_{p_{\theta,t}}(\bm{x}_{t})-\bm{s}_{q_{t}}(\bm{x}_{t}))\bigg{\}}\mathrm{d}t,$ (3.2)
where $p_{\theta,t}$ and $q_{t}$ denote the marginal densities of the diffusion process (2.1) at time $t$ initialized with $p_{\theta,0}=p_{\theta}$ and $q_{0}=p_{ref}$ respectively. $w(t)$ is an integral weighting function. Clearly, we have $\mathbf{D}^{[0,T]}(p_{\theta},p_{ref})=0$ if and only if $p_{\theta}(\bm{x}_{0})=p_{ref}(\bm{x}_{0}),~{}a.s.~{}\pi_{0}$.
3.2 Diff-Instruct*
Recall that $g_{\theta}$ is a one-step model, therefore samples from $p_{\theta}$ can be implemented through a direct mapping $\bm{x}_{0}=g_{\theta}(\bm{z},\bm{c})$. With the score-based regularization term (3.2), for each given text prompt $\bm{c}$, we can formally write down our training objective to minimize as:
$\mathcal{L}_{Orig}(\theta)=\mathbb{E}_{\bm{z}\sim p_{z},\atop\bm{x}_{0}=g_{\theta}(\bm{z},\bm{c})}\big{[}-\alpha r(\bm{x}_{0},\bm{c})\big{]}+\mathbf{D}^{[0,T]}(p_{\theta},p_{ref})$ (3.3)
Now we are ready to reveal the objective of Diff-Instruct* that we use to train human-preferred one-step models $g_{\theta}$. Notice that directly minimizing objective (3.3) is intractable because we do not know the relationship between $\theta$ and corresponding $p_{\theta,t}$. However, we show in Theorem 3.1 that an equivalent tractable loss (3.4) will have the same $\theta$ gradient as the intractable loss function (3.3):
$\mathcal{L}_{DI*}(\theta)=$ $\mathbb{E}_{\bm{z}\sim p_{z},\atop\bm{x}_{0}=g_{\theta}(\bm{z})}\bigg{[}-\alpha r(\bm{x}_{0},\bm{c})$ (3.4)
$+\int_{t=0}^{T}w(t)\mathbb{E}_{\bm{x}_{t}|\bm{x}_{0}\atop\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})}\bigg{\{}-\mathbf{d}^{\prime}(\bm{y}_{t})\bigg{\}}^{T}\bigg{\{}\bm{s}_{p_{\operatorname{sg}[\theta],t}}(\bm{x}_{t})-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\bigg{\}}\mathrm{d}t\bigg{]}$
with $\bm{y}_{t}\coloneqq\bm{s}_{p_{\operatorname{sg}[\theta],t}}(\bm{x}_{t})-\bm{s}_{q_{t}}(\bm{x}_{t})$.
Theorem 3.1.
Under mild assumptions, if we take the sampling distribution in (3.2) as $\pi_{t}=p_{\operatorname{sg}[\theta],t}$, then the $\theta$ gradient of (3.3) is the same as the objective (3.4): $\frac{\partial}{\partial\theta}\mathcal{L}_{Orig}(\theta)=\frac{\partial}{\partial\theta}\mathcal{L}_{DI*}(\theta)$.
In practice, we can use another assistant diffusion model $\bm{s}_{\psi}(\bm{x}_{t},t)$ to approximate the 1-step model’s score function $\bm{s}_{p_{\operatorname{sg}[\theta],t}}(\bm{x}_{t})$ pointwise, which was also done in the literature of diffusion distillations works such as Luo et al. (2024b); Zhou et al. (2024a; b); Luo et al. (2024c); Yin et al. (2023; 2024). Therefore, we can alternate between 1) updating the assistant diffusion $\bm{s}_{\psi}(\bm{x}_{t},t)$ using 1-step model generated samples (which are efficient) and 2) updating the 1-step model by minimizing the tractable objective (3.4). We name our training method that minimizes the objective $\mathcal{L}_{DI*}(\theta)$ in (3.4) the Diff-Instruct* because it is inspired by Diff-Instruct(Luo et al., 2024b) and Diff-Instruct++(Luo, 2024) that involves an additional diffusion model and a reward model to train one-step t2i models.
3.3 Two Ways of Incorporating Classifier-free Guidance
Input: prompt dataset $\mathcal{C}$, 1-step model $g_{\theta}(\bm{x}_{0}|\bm{z},\bm{c})$, prior distribution $p_{z}$, reward model $r(\bm{x},\bm{c})$, reward model scale $\alpha_{rew}$, CFG reward scale $\alpha_{cfg}$, reference diffusion model $\bm{s}_{ref}(\bm{x}_{t}|c,\bm{c})$, assistant diffusion $\bm{s}_{\psi}(\bm{x}_{t}|t,\bm{c})$, forward diffusion $p_{t}(\bm{x}_{t}|\bm{x}_{0})$ (2.1), assistant diffusion updates rounds $K_{TA}$, time distribution $\pi(t)$, diffusion model weighting $\lambda(t)$, student loss time weighting $w(t)$.
while not converge do
       freeze $\theta$, update ${\psi}$ for $K_{TA}$ rounds using SGD by minimizing
$\mathcal{L}(\psi)=\mathbb{E}_{\bm{c}\sim\mathcal{C},\bm{z}\sim p_{z},t\sim\pi(t)\atop\bm{x}_{0}=g_{\theta}(\bm{z}|\bm{c}),\bm{x}_{t}|\bm{x}_{0}\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})}\lambda(t)\|\bm{s}_{\psi}(\bm{x}_{t}|t,\bm{c})-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\|_{2}^{2}\mathrm{d}t.$
freeze $\psi$, update $\theta$ using SGD by minimizing loss
$\mathcal{L}_{DI*}(\theta)=$ $\mathbb{E}_{\bm{c}\sim\mathcal{C},\bm{z}\sim p_{z},\atop\bm{x}_{0}=g_{\theta}(\bm{z},\bm{c})}\bigg{\{}-\alpha_{rew}\cdot r(\bm{x}_{0},\bm{c})+\mathbb{E}_{t\sim\pi(t),\atop\bm{x}_{t}|\bm{x}_{0}\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})}\bigg{[}$
$-w(t)\big{\{}\mathbf{d}^{\prime}(\bm{s}_{\psi}(\bm{x}_{t}|t,\bm{c})-\bm{s}_{ref}(\bm{x}_{t}|t,\bm{c}))\big{\}}^{T}\big{\{}\bm{s}_{\psi}(\bm{x}_{t}|t,\bm{c})-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\big{\}}$
$+\alpha_{cfg}\cdot w(t)\big{\{}\bm{s}_{ref}(\operatorname{sg}[\bm{x}_{t}]|t,\bm{c})-\bm{s}_{ref}(\operatorname{sg}[\bm{x}_{t}]|t,\bm{\varnothing})\big{\}}^{T}\bm{x}_{t}\bigg{]}\bigg{\}}$ (3.5)
end while
return $\theta,\psi$.
Algorithm 1 Diff-Instruct* for training human-preferred one-step text-to-image 1-step model.
Classifier-free Guidance Secretly Induced an Implicit Reward.
In previous sections, we have shown in theory that with explicitly available reward models, we can readily train the one-step models to align with human preference. In this section, we enhance the DI* by incorporating the classifier-free reward that is implied by the classifier-free guidance of diffusion models.
The classifier-free guidance (Ho & Salimans, 2022) (CFG) uses a modified score function of a form
$\tilde{\bm{s}}_{ref}(\bm{x}_{t},t|\bm{c})\coloneqq\bm{s}_{ref}(\bm{x}_{t},t|\bm{\varnothing})+\omega\big{\{}\bm{s}_{ref}(\bm{x}_{t},t|\bm{c})-\bm{s}_{ref}(\bm{x}_{t},t|\bm{\varnothing})\big{\}}$
to replace the original conditions score function $\bm{s}_{ref}(\bm{x}_{t},t|\bm{c})$. Using CFG for diffusion models empirically leads to better sampling quality.
As is first pointed out by Luo (2024), the classifier-free guidance is related to an implicit reward function. In this part, we derive a tractable loss function that minimizes the so-called classifier-free reward, which we use together with the explicit reward $r(\cdot,\cdot)$ in DI*.
Theorem 3.2.
Under mild conditions, if we set an implicit reward function as (3.7), the loss (3.6)
$\mathcal{L}_{cfg}(\theta)=\int_{t=0}^{T}\mathbb{E}_{\bm{z}\sim p_{z},\bm{x}_{0}=g_{\theta}(\bm{z},\bm{c})\atop\bm{x}_{t}|\bm{x}_{0}\sim p(\bm{x}_{t}|\bm{x}_{0})}w(t)\bigg{\{}\bm{s}_{ref}(\operatorname{sg}[\bm{x}_{t}]|t,\bm{c})-\bm{s}_{ref}(\operatorname{sg}[\bm{x}_{t}]|t,\bm{\varnothing})\bigg{\}}^{T}\bm{x}_{t}\mathrm{d}t$ (3.6)
has the same gradient as the negative implicit reward function (3.7)
$-r(\bm{x}_{0},\bm{c})=-\int_{t=0}^{T}\mathbb{E}_{\bm{x}_{t}\sim p_{\theta,t}}w(t)\log\frac{p_{ref}(\bm{x}_{t}|t,\bm{c})}{p_{ref}(\bm{x}_{t}|t)}\mathrm{d}t.$ (3.7)
The notation $\operatorname{sg}[\bm{x}_{t}]$ means detaching the $\theta$ gradient on $\bm{x}_{t}$. Theorem 3.2 gives a tractable loss function (3.6) aiming to minimize the negative classifier-free reward function. Therefore, we can scale and add this loss $\mathcal{L}_{cfg}(\theta)$ (3.6) to the DI* loss (3.4) to balance the effects of explicit reward and implicit CFG reward.
Figure 3: A visual comparison on Parti prompts of our introduced 2.6B DI*-SDXL-1step text-to-image generative model and 12B FLUX-dev-50step and SD3.5-large-28step diffusion models. DI*-SDXL-1step model outperforms FLUX-dev-50step significantly in PickScore, ImageReward, and CLIPScore, while only taking 1.88% inference time and 29.30% GPU memory cost. Benefiting from RLHF using human feedback reward, DI*-SDXL-1step shows better aesthetic details, improved layout, and better reality than FLUX-dev-50step and SD3.5-large-28step. Prompts from top to bottom are “the skyline of New York City”, “a shiba inu”, “Pirate ship sailing into a bioluminescence sea with a galaxy in the sky, epic, 4k, ultra”, and “a volcano exploding next to a marina”.
Choice 1: Incorporating CFG via Implicit CFG Reward. Based on Theorem 3.2, we can incorporate classifier-free guidance by adding the pseudo reward (3.7) to explicit human reward to get a mixed reward. The guidance scale $\alpha_{cfg}$ and $\alpha_{rew}$ balances the strength of explicit human reward and implicit CFG reward. Choice 2: Incorporate CFG via CFG-enhanced Reference Diffusion. Besides the implicit CFG reward, we can also inject the CFG mechanism into Diff-Instruct* by replacing the naive reference diffusion with a CFG-enhanced reference diffusion which writes $\tilde{\bm{s}}_{ref}(\bm{x}_{t}|t,\bm{c})\coloneqq\bm{s}_{ref}(\bm{x}_{t}|t,\bm{c})+\alpha_{cfg}(\bm{s}_{ref}(\bm{x}_{t}|t,\bm{\varnothing}))$. This approach is straightforward. In practice, we find empirically that the choice 2 is easier to tune and results in better performances. However, choice 1 also leads to solid 1-step models. Please check out ablation studies in Table 2 for details.
3.4 The Practical Algorithm
Now it is time for us to introduce the practical algorithm. As Algorithm 1 shows, the DI* involves three models, with one 1-step model $g_{\theta}$, one reference diffusion model $\bm{s}_{ref}$ and one assistant diffusion model $\bm{s}_{\psi}$. The reference diffusion does not need to be trained, while the 1-step model and the assistant diffusion are updated alternatively. Two hyper-parameters, the $\alpha_{rew}$ and $\alpha_{cfg}$ control the strength of the explicit reward and the implicit CFG reward during training. The explicit reward model can either be an off-the-shelf reward model, such as PickScore (Kirstain et al., 2023), CLIP similarity score (Radford et al., 2021), or reward trained in-house with internal feedback data.
Pseudo-Huber Distance Functions.
Clearly, various choices of distance function $\mathbf{d}(.)$ result in different training algorithms. For instance, $\mathbf{d}(\bm{y}_{t})=\|\bm{y}_{t}\|_{2}^{2}$ is a naive choice. Interestingly, such a distance function has been studied in pure diffusion distillation literature in Luo et al. (2024c); Zhou et al. (2024b; a). In this paper, we draw inspiration from (Luo et al., 2024c) and find that using the so-called pseudo-Huber distance leads to better performance. The distance and corresponding loss writes $\bm{d}(\bm{y})\coloneqq\sqrt{\|\bm{y}_{t}\|_{2}^{2}+c^{2}}-c$, and $\bm{y}_{t}\coloneqq\bm{s}_{p_{\operatorname{sg}[\theta],t}}(\bm{x}_{t})-\bm{s}_{q_{t}}(\bm{x}_{t})$, then
$\mathbf{D}^{[0,T]}(p_{\theta},p_{ref})=-\bigg{\{}\frac{\bm{y}_{t}}{\sqrt{\|\bm{y}_{t}\|_{2}^{2}+c^{2}}}\bigg{\}}^{T}\bigg{\{}\bm{s}_{\psi}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\bigg{\}}.$ (3.8)
Figure 4: Ablation comparisons of DI*-1step models and other SDXL-based models in Table 2.
4 Related Works
Diffusion Distillation Through Divergence Minimization.
Diff-Instruct* is inspired by research on diffusion distillation (Luo, 2023) which aims to minimize distribution divergence to train one-step text-to-image models. Luo et al. (2024b) first study the diffusion distillation by minimizing the Integral KL divergence. Yin et al. (2023) generalize Diff-Instruct and add a data regression loss for better performance. Zhou et al. (2024b) study the distillation by minimizing the Fisher divergence. Luo et al. (2024c) study the distillation using the general score-based divergence. Many other works also introduced additional techniques and improved the performance (Geng et al., 2023; Kim et al., 2023; Song et al., 2023; Song & Dhariwal, ; Nguyen & Tran, 2023; Song et al., 2024; Yin et al., 2024; Zhou et al., 2024a; Heek et al., 2024; Xie et al., 2024; Salimans et al., 2024; Geng et al., 2024; Meng et al., 2022; Sauer et al., 2023b; Luo et al., 2023a; Liu et al., 2023; Gu et al., 2023; Xu et al., 2024; Ren et al., 2024; Lin et al., 2024; Zheng et al., 2024; Li et al., 2024; Berthelot et al., 2023; Zheng et al., 2022).
Preference Alignment for Diffusion Models and One-step Models.
In recent years, many works have emerged trying to align diffusion models with human preferences. There are three main lines of alignment methods for diffusion models. 1) The first kind of method fine-tunes the diffusion model over a specifically curated image-prompt dataset (Dai et al., 2023; Podell et al., 2023). 2) the second line of methods tries to maximize some reward functions either through the multi-step diffusion generation output (Prabhudesai et al., 2023; Clark et al., 2023; Lee et al., 2023) or through policy gradient-based RL approaches (Fan et al., 2024; Black et al., 2023). For these methods, the backpropagation through the multi-step diffusion generation output is expensive and hard to scale. 3) the third line, such as Diffusion-DPO (Wallace et al., 2024), Diffusion-KTO (Yang et al., 2024), tries to directly improve the diffusion model’s human preference property with raw collected data instead of reward functions. Besides the human preference alignment of diffusion models, Diff-Instruct++(Luo, 2024) recently arose as the first attempt to improve human preferences for one-step text-to-image models. Though we get inspiration from Diff-Instruct++, our Diff-Instruct* uses score-based divergences which are technically different from the KL divergence used in DI++. Besides, as we show in Section 5.1, Diff-Instruct* archives better performances than Diff-Instruct++.
Table 1: Quantitative comparisons of $1024\times 1024$ resolution leading text-to-image models on Parti(Yu et al., 2022) Prompts (the upper part) and MSCOCO-2017 validation prompts (the under part). DI*: short for Diff-Instruct*. SIM: short for Score Implicit Matching. : our implementation.
[…]
Parti Prompts
SDXL-Base(Podell et al., 2023) 50 UNet 2.6B 0.887 5.72 0.2274 32.72 111 Sec
SDXL-DPO(Wallace et al., 2024) 50 UNet 2.6B 1.102 5.77 0.2290 33.03 111 Sec
SD3.5-large(sd, 3) 28 DiT 8B 1.133 5.70 0.2306 32.70 66.23 Sec
FLUX-dev(flu, ) 50 DiT 12B 1.132 5.90 0.2317 31.70 118.64 Sec
DMD2-SDXL(Yin et al., 2024) 1 UNet 2.6B 0.930 5.51 0.2249 32.97 2.22 Sec
Diff-Instruct(Luo et al., 2024b) 1 UNet 2.6B 1.058 5.60 0.2253 33.02 2.22 Sec
SIM (Luo et al., 2024c) 1 UNet 2.6B 1.049 5.66 0.2273 32.93 2.22 Sec
Diff-Instruct++-SDXL(Luo, 2024) 1 UNet 2.6B 1.061 5.58 0.2260 32.94 2.22 Sec
DI*-SDXL (Ours) 1 UNet 2.6B 1.067 5.74 0.2304 32.82 2.22 Sec
DI*-SDXL (Ours, Longer Training) 1 UNet 2.6B 1.140 5.83 0.2331 32.75 2.22 Sec
[…]
SDXL-Base(Podell et al., 2023) 50 UNet 2.6B 0.825 5.55 0.2281 31.86 15.89
SDXL-DPO(Wallace et al., 2024) 50 UNet 2.6B 0.950 5.69 0.2295 32.17 21.48
SD3.5-large(sd, 3) 28 DiT 8B 1.048 5.46 0.2305 31.85 17.69
FLUX-dev(flu, ) 50 DiT 12B 1.065 5.64 0.2333 30.88 24.77
DMD2-SDXL(Yin et al., 2024) 1 UNet 2.6B 0.822 5.42 0.2251 31.89 15.80
Diff-Instruct-SDXL(Luo et al., 2024b) 1 UNet 2.6B 0.907 5.48 0.2264 31.85 21.62
SIM-SDXL(Luo et al., 2024c) 1 UNet 2.6B 0.925 5.54 0.2277 31.90 18.84
Diff-Instruct++-SDXL(Luo, 2024) 1 UNet 2.6B 0.908 5.53 0.2265 31.85 21.42
DI*-SDXL (Ours) 1 UNet 2.6B 0.933 5.57 0.2305 31.75 19.09
DI*-SDXL (Ours, Longer Training) 1 UNet 2.6B 0.983 5.65 0.2335 31.57 21.15
Table 2: Ablation study on Parti(Yu et al., 2022) Prompts of 1024 resolution DI*-SDXL-1step models with different reward scales. DI*: short for Diff-Instruct*. SIM: short for Score Implicit Matching. DI*-Out: incorporating CFG with implicit CFG reward; DI*-In: incorporating CFG with enhanced reference diffusion. : our implementation.
[…]
DMD2-SDXL(Init Model) 1 2.6B 0.938 5.51 0.2249 32.97 --
DI++-SDXL (Aligned using KL) 1 2.6B 0.846 5.50 0.2243 32.66 $(0,0)$
DI++-SDXL (equ to Diff-Instruct) 1 2.6B 1.058 5.60 0.2253 33.02 $(0,7.5)$
DI++-SDXL (Aligned using KL) 1 2.6B 1.061 5.58 0.2260 32.94 $(100,7.5)$
DI*-Out-SDXL (Out CFG) 1 2.6B 1.082 5.63 0.2263 33.03 $(100,7.5)$
DI*-In-SDXL ( Baseline, No Reward) 1 2.6B 0.782 5.74 0.2256 32.16 $(0,0)$
DI*-In-SDXL (equ to SIM, Only CFG) 1 2.6B 1.049 5.66 0.2273 32.93 $(0,7.5)$
DI*-In-SDXL ((Human Reward + CFG)) 1 2.6B 1.031 5.69 0.2274 32.87 $(1,7.5)$
DI*-In-SDXL 1 2.6B 1.048 5.66 0.2278 32.91 $(10,7.5)$
DI*-In-SDXL 1 2.6B 1.020 5.68 0.2278 32.82 $(100,4.5)$
DI*-In-SDXL 1 2.6B 1.067 5.74 0.2304 32.82 $(100,7.5)$
DI*-In-SDXL (Best, longer training) 1 2.6B 1.140 5.83 0.2331 32.75 $(100,7.5)$
5 Experiments
Experiment Settings for SDXL Experiment.
We use the open-sourced SDXL of 1024$\times$1024 resolution as our reference diffusion in Algorithm 1. We construct the 1-step model with the same architecture as the SDXL model and initialize the 1-step model weights with DMD2-SDXL-1step model (Yin et al., 2024), which is a good 1-step diffusion distillation model using a combination of KL divergence and GAN loss. We use the prompts of the LAION-AESTHETIC dataset with an aesthetic score larger than 6.25, which resulting a total of 3M text prompts. We use the PickScore (Kirstain et al., 2023)(a high-quality off-the-shelf reward model trained using human feedback data) as the explicit human reward. In Table 2, we do an ablation study that compares 1-step DI* models with different guidance strategies and reward scales and finds that using inner guidance with a scale of 7.5 and an explicit human reward with a scale of 100 results in the best performances.
Quantitative Evaluations Metrics.
We compare the DI*-SDXL-1step model with other leading open-sourced models of a resolution of $1024\times 1024$ on the Parti prompt benchmark, COCO-2014-30K benchmark, and also Human preference scores. These models include diffusion-based (or flow-matching) models such as FLUX-dev (flu, ), Stable-Diffusion-3.5-Large(sd, 3), and SDXL (Podell et al., 2023) with its DPO variant (Wallace et al., 2024); 1024 resolution 1-step model such as DMD2(Yin et al., 2024), Diff-Instruct(Luo et al., 2024b), Score Implicit Matching (SIM) (Luo et al., 2024c), and also Diff-Instruct++ (Luo, 2024). On Parti and COCO prompt datasets, we compute four standard scores in Table 1: the Image Reward (Xu et al., 2023), the Aesthetic Score (Schuhmann, 2022), the PickScore(Kirstain et al., 2023), and the CLIP score(Radford et al., 2021). We also evaluate the 30K-FID (Heusel et al., 2017) on the COCO validation dataset. On Human Preference Score (Wu et al., 2023) benchmarks, we calculate HPSv2.1 scores.
Table 3: Quantitative evaluations of models on HPSv2.1 scores. We compare open-sourced models regardless of their base model and architecture. $\dagger$ indicates our implementation.
[…]
50Step-SDXL-base(Podell et al., 2023) $30.85$ $29.30$ $28.98$ $27.05$ $29.05$
50Step-SDXL-DPO(Wallace et al., 2024) $32.01$ $30.75$ $30.70$ $28.24$ $30.42$
28Step-SD3.5-large $31.89$ $30.19$ $30.39$ $28.01$ $30.12$
50Step-FLUX-dev $32.09$ $30.44$ $31.17$ 29.09 $30.70$
1Step-DMD2-SDXL(Yin et al., 2024) $29.72$ $27.96$ $27.64$ $26.55$ $27.97$
1Step-Diff-Instruct-SDXL(Luo et al., 2024b) $31.15$ $29.71$ $29.72$ $28.20$ $29.70$
1Step-SIM-SDXL(Luo et al., 2024c) $31.97$ $30.46$ $30.13$ $28.08$ $30.16$
1Step-DI++-SDXL(Luo, 2024) $31.19$ $29.88$ $29.61$ $28.21$ $29.72$
1Step-DI*-SDXL(ours) $32.26$ $30.57$ $30.10$ $27.95$ $30.22$
1Step-DI*-SDXL(ours, longer training) 33.22 31.67 31.25 28.6228.62{28.62}28.62 31.19
5.1 Performances and Findings
2.6B DI*-SDXL-1step Model Shows Very Strong Performances.
The FLUX-dev-50step model(flu, ) is regarded as a leading open-sourced diffusion model that is trained over millions of high-quality internal data. As Table 1 shows, our best 2.6B DI*-SDXL-1step model slightly outperforms the 12B FLUX-dev-50step diffusion model and 8B SD3.5-large-28step with only 1.88% inference time and 29.3% GPU memory costs on the Parti prompt dataset which consists of challenging prompts with rich concepts. The 2.6B DI*-SDXL-1step is on par with FLUX and SD3.5 on COCO prompts in terms of preference scores and FIDs. As Table 3 shows, the DI*-SDXL-1step model has a record-breaking HPSv2.1 score of 31.19, outperforming the rest of text-to-image models of $1024\times 1024$ resolution.
Ablation Study.
In Table 2, we conduct a comprehensive ablation study on the Parti prompt dataset to compare DI* (RLHF using score-based divergence) with different guidance strategies and different reward scales (which sometimes recover previous distillation approaches such as Diff-Instruct(Luo et al., 2024b) and Score Implicit Matching(Luo et al., 2024c)), and DI++(Luo, 2024)(RLHF using KL divergence). The results show that DI* using score-based divergence outperforms the DI++ model under the same configurations and reward scales. Besides, with proper reward scales, DI* models show dominant human preference scores, indicating its simple yet effective contribution to aligning one-step text-to-image models with human feedback. Besides, we also find that DI++ using KL divergence is very easy to collapse to painting-like images with over-saturated color and lighting (please refer to Figure 1 for visualizations) that have high human reward but lack diversities. Instead, using score-based divergences empirically maintains diversity while still steadily improving human preferences. This means DI* can improve images’ human preference regardless they are painting-like, photo-like, or animations. However, we do acknowledge an interesting finding of the trade-off between preference scores and CLIP scores: we find that the larger-scale explicit human reward is assigned, the worse the CLIPS score will be. We think this phenomenon reveals a natural contradiction between human-preferred data distributions and objective data distribution. The RLHF encourages the 1-step model to generate human-preferred images, which inevitably changes the underlying distribution from the ground-truth data distribution. We also find that though FLUX shows very strong human preference performances, its objective scores such as the CLIPScore and the FID are not that good even compared with smaller models.
Qualitative Comparisons.
Figure 2 shows some 1024$\times$1024 images generated by DI*-SDXL-1step models. Figure 3 shows a qualitative comparison of the DI*-SDXL-1step model against FLUX and SD3.5 diffusion models. We are surprised to find that the DI*-SDXL-1step model shows better aesthetic details, improved layouts, and reality lighting than the other two models due to the RLHF. Figure 4 shows the visual comparison of DI* (with different ablative reward scales) against our initial model (the DMD2-1step model(Yin et al., 2024)), other 1-step models, and SDXL with and without DPO. While SIM (Luo et al., 2024c), DI++(Luo, 2024), and Diff-Instruct(Luo et al., 2024b) can output high-quality samples, they seem to suffer from over-saturated issues, meaning that the images have over-saturated colors and lighting that might make users feel uncomfortable. On the contrary, the DI*-1step model shows very gentle lighting and aesthetic colors, which align better with human preferences. We empirically conclude that DI* prevents the model from collapsing into high-reward painting-like images in which DI++ collapsed. Besides, the SDXL and SDXL-DPO are easy to generate painting-like images instead of photo-realistic images.
6 Conclusion and Limitations
In this paper, we present Diff-Instruct*, a novel approach for preference alignment of 1-step text-to-image generative models. We also introduced two novel approaches to incorporate traditional classifier-free guidance into alignment through the lens of explicit-implicit reward decoupling. We Diff-Instruct*, we introduce the 2.6B DI*-SDXL-1step model that outperforms current leading models such as the 12B FLUX-dev-50step model with merely 1.88% inference time cost.
Nonetheless, Diff-Instruct* has its limitations. We empirically find some typical mistakes the 1-step model often makes. For instance, the 1-step model Sometimes Generates Bad Human Faces and Hands. We believe that consistently improving the 1-step model architecture, scaling its parameters, improving reference diffusions, and enhancing the reward models can lead to better models. Besides, we are also interested in following directions that call for further research. For instance, the DI* needs a pre-trained reward model. However, methods like DPO (Rafailov et al., 2024; Wallace et al., 2024) have put a new setup that trains generative models directly using human feedback data. We hope these future directions could contribute more to the community.